# A Holistic Optimization Framework for Energy Efficient UAV-assisted Fog Computing: Attitude Control, Trajectory Planning and Task Assignment

### v2

@@ V Algorithms
V-A PSO for Task Assignment and Resource Allocation
We first introduce an algorithm based on the idea of Particle Swarm Optimization (PSO), for assigning tasks and allocating resources given that the trajectory is already known. The steps of our proposed algorithm are demonstrated in Algorithm 1, which we refer the readers to the conference version of this paper [15] for more detailed explanations. We categorize the job assignment decision, transmission power, and processing frequency as a collective entity $s_{h}=(\mathbf{o}_{ij},p_{j}(t),f_{j}(t))$ within the particle swarm. Here, $h\in\{1,...,H\}$. Next, the determination of the particle group’s position is carried out by uniformly sampling a set of particle groups denoted by $H\in\mathbb{N_{+}}$. In order to search for the optimal solution, the velocity of the particle groups is initialized and subsequently updated using the method in [47]. The algorithm terminates when the discrepancy between the outcomes of $20$ successive rounds is less than a specified threshold $\eth$, and then outputs the minimal value $\mathbb{S}^{*}$, the optimal decision for job assignment $\mathbf{o}_{ij}^{*}$, the transmission power $p_{j}^{*}(t)$, and the processing frequency $f_{j}^{*}(t)$.
@@ V-A PSO for Task Assignment and Resource Allocation
4:     Initialize particles’ position $G_{h}$
5:     Initialize particles’ velocity $V_{h}$
@@
8:for $h=1$ to $H$ do
9:     Update $V_{h}$ and $G_{h}$ by acceleration factors $F_{A1}$, $F_{A2}$ and $F_{I}$
10:     if $fit(G_{h})<fit(pBest_{h})$ then
11:         $pBest_{h}\longleftarrow G_{h}$
12:         if $fit(pBest_{h})<fit(gBest)$ then
13:              $gBest(k)\longleftarrow pBest_{h}$
14:     $k\longleftarrow(k+1)$
15:while $\lvert gBest(k+1)-gBest(k)\rvert<\eth$ do
16:     $\mathbb{S}^{*}(t)\longleftarrow fit(gBest)$
17:     $(\mathbf{o}_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t))$ $\longleftarrow gBest$
18:Output $\mathbb{S}^{*}(t)$; $(\mathbf{o}_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t))$
V-B ACS-DS for Trajectory Planning: Motivation and Concept
We now present the concept and outlines of the ACS-DS (ACS with Decoupling and Safety values) algorithm to generate the optimal UAV trajectory to minimize $E^{\mathrm{MOV}}$. Then, we will prove that ACS-DS always converges to the optimal solution in polynomial time in subsequent subsections.
@@ V-B2 Decoupling
V-C ACS-DS: Detailed Steps
We let $m$ be the total number of ants in the colony, and $\sigma=V_{P}+\kappa$ be the guidance factor. We further denote $\sigma_{\mu\nu}(h)$ as the sum of pheromone values and safety values between the neighboring waypoints $\mu$ and $\nu$ in the $h$th iteration. The initial pheromones on each edge are equal, namely $\sigma_{\mu\nu}(0)=C$ for all $\mu$ and $\nu$. We denote by $s^{*}$ the global best path obtained by the algorithm, corresponding to the minimal-cost trajectory in our UAV optimization problem. Next, for each ant $k\in\{1,2\ldots,m\}$ in the colony, we initialize the pheromone $V_{P0}$, the safety value $\kappa_{0}$, and the heuristic value $\eta$. We also define the evaporation rate $\rho\in(0,1)$ representing the degree to which the guidance factor $\sigma_{\mu\nu}(h)$ decays with iterations. Finally, we define $S_{\iota}$ as the set of waypoints that ant $\iota$ can pass next, and $p_{\mu\nu}^{\iota}(h)$ as the probability that ant $\iota$ moves from position $\mu$ to $\nu$ in the $h$th iteration. The exact value of $p_{\mu\nu}^{\iota}(h)$ is jointly determined by the guidance factor and heuristic value at the waypoint, as in the following equation,
$p_{\mu\nu}^{\iota}(h)=\begin{cases}\frac{\sigma_{\mu\nu}^{\alpha}(h)\eta_{\mu\nu}{}^{\beta}(h)}{\sum_{r\in S_{\iota}}\sigma_{ir}^{\alpha}\eta_{ir}{}^{\beta}(h)}&\nu\in S_{\iota},\\ \quad\quad\quad 0&\text{ otherwise.}\end{cases}$ (23)
@@ V-C ACS-DS: Detailed Steps
1:Positions of all MDs $(x_{j},y_{j},z_{j})$; $\rho$; $\eta$; $V_{P0}$; $\kappa_{0}$; terrain surface with no-fly zone $(x^{o}_{n},y^{o}_{n},z^{o}_{n})\in\boldsymbol{O}$
2:$E^{\mathrm{MOV}*}(t)$; Trajectory of the UAV $(x_{\iota}(t),y_{\iota}(t),z_{\iota}(t))$
@@
8:         for each ant $\iota$ do
9:              $(x_{\iota}(h),y_{\iota}(h),z_{\iota}(h))$ = $(x(0),y(0),z(0))$
10:              for each edge $(\mu,\nu)$ do
11:                  Choose the next coordinate with probability $p_{\mu\nu}^{\iota}(h)$ by $\sigma$ and $\eta$
12:                  while $(x_{\iota}(h+1),y_{\iota}(h+1),z_{\iota}(h+1))\notin\boldsymbol{O}$ do
13:                       Output $(x_{\iota}(h+1),y_{\iota}(h+1),z_{\iota}(h+1))$
14:              Compute and output the length $\sum_{t=1}^{T}d_{u}(t)$ of the path by the $\iota$th ant and $E_{u}$
15:              for each edge $(\mu,\nu)$ do
16:                  Update $V_{P_{\mu\nu}}$ and $\kappa_{\mu\nu}$ by $\rho$
17:                  Update $\sigma_{\mu\nu}(h)$ by $V_{P_{\mu\nu}}$ and $\kappa_{\mu\nu}$
18:Compute and output $E^{\mathrm{MOV}*}(t)$ by (10).
V-D ACS-DS: Proof of convergence
@@
$\lim_{h\rightarrow\infty}\sigma_{\mu^{\prime}\nu^{\prime}}\left(h^{*}+h^{\prime}\right)$ $=\max\lim_{h\rightarrow\infty}\left\{\sigma_{\mu\nu}\left(h^{*}\right),\left[(1-\rho)^{h^{\prime}}\cdot\sigma_{\mu^{\prime}\nu^{\prime}}\left(h^{*}\right)\right]\right\}$ (26)
@@
$\leq\sigma_{\mu\nu}\left(h^{*}\right)$
@@ V-E ACS-DS: Complexity analysis
VI Performance Evaluation
In this section, we present numerical results of simulations on systems with various settings to assess the efficacy and versatility of our proposed solutions.
VI-A Experiment setup
We consider a three-dimensional spatial domain with dimensions of $S\times S\times Z$. $K$ MDs are uniformly and randomly distributed in the space. Values of key system parameters are listed in Table V.
TABLE V: Values of relevant system parameters in the experiments
@@ VI-A Experiment setup
$m$ 80 $\mathrm{~{}kg}$ $a$ 5.28 $\mathrm{~{}m}/\mathrm{s}^{2}$
@@
The values of parameters used in the algorithms described earlier in this paper are listed in Table VI. All experimental results presented in this section are based on the average of the 50 independent runs with the weighting factor $\epsilon$ randomly generated randomly within its domain for each run.
TABLE VI: Values of relevant algorithmic parameters
@@
For the task assignment and resource allocation part, we adopt PSO (Algorithm 1), which has been demonstrated effective and robust in our earlier conference paper [15]), for all experiments in this section. For the attitude control and trajectory planning modules, we focus on evaluating the performance of our proposed ACS-DS (Algorithm 2) and ATC (Section IV-A). Specifically, we examine the following eight implementations and compare their performances in terms of overall operational efficiency cost.
  • RAN: All tasks are randomly assigned to the MD, the UAV, or the data center. The channels, processing frequencies and power allocations are generated randomly in their respective domains. No special mechanisms are applied for attitude control or trajectory planning.
@@
  • PSO: The PSO algorithm (Algorithm 1) is used to determine task assignments, the processing frequencies, and transmission powers of MDs and the UAV. Allocation of channels in accordance with the methodology described in Section III. No special mechanisms are applied for attitude control or trajectory planning.
  • ACS: On top of PSO, trajectory of the UAV is planned by the classical ACS.
  • ACS $+$ ATC: On top of ACS, the mechanism described in Section IV-A is added to control the attitude of the UAV.
  • ACS-D: On top of PSO, trajectory of the UAV is planned by ACS with the decoupling mechanism.
  • ACS-D $+$ ATC: The mechanism described in Section IV-A is added to control the attitude of the UAV on top of ACS-D.
  • ACS-DS: On top of PSO. Trajectory of the UAV is planned by ACS with decoupling and safety values mechanisms in Algorithm 2.
  • ACS-DS $+$ ATC: The mechanism described in Section IV-A is added to control the attitude of the UAV on top of ACS-DS.
VI-B Numerical results
VI-B1 Optimal Propeller Parameters

### v3

@@ 5.2 Task Assignment and Resource Allocation
[…]
Regarding fairness considerations, we note that fairness is not explicitly included as an optimization target in our objective function, which focuses on minimizing overall energy consumption and delay. However, our system design incorporates natural fairness mechanisms through orthogonal channel allocation (preventing channel monopolization) and the feasibility constraints that ensure all tasks are assigned to appropriate execution locations. The Gamma-based channel allocation strategy further promotes equitable resource distribution by preventing extremely large tasks from dominating available channels. A comprehensive quantitative fairness analysis, including metrics such as Jain’s fairness index or per-user delay variance, would require additional modeling extensions and is considered as future work for scenarios where explicit fairness guarantees are required.
@@
4:  Initialize particles’ positions $G_{h}$
5:  Initialize particles’ velocities $V_{h}$
@@
9:for $h=1$ to $H$ do
10:  Update $V_{h}$ and $G_{h}$ by acceleration factors $F_{A1}$, $F_{A2}$ and $F_{I}$
11:  if $fit(G_{h})<fit(pBest_{h})$ then
12:   $pBest_{h}\longleftarrow G_{h}$
13:   if $fit(pBest_{h})<fit(gBest)$ then
14:     $gBest(k)\longleftarrow pBest_{h}$
15:   end if
16:  end if
17:  $k\longleftarrow(k+1)$
18:end for
[…]
20:  $\mathbb{S}^{*}(t)\longleftarrow fit(gBest)$
21:  $(\mathbf{o}_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t))$ $\longleftarrow gBest$
22:end while
23:Output $\mathbb{S}^{*}(t)$; $(\mathbf{o}_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t))$
5.3 ACS-DS for Trajectory Planning: Motivation and Concept
We now present the concept and outlines of the ACS-DS (Ant Colony System with Decoupling and Safety values) algorithm to generate the optimal UAV trajectory to minimize $E^{\mathrm{MOV}}$. Then, we will prove that ACS-DS always converges to the optimal solution in polynomial time in subsequent subsections.
@@ 5.3.2 Decoupling
5.4 ACS-DS: Detailed Steps
We let $m$ be the total number of ants in the colony, and $\sigma=V_{P}+\kappa$ be the guidance factor. We further denote $\sigma_{\mu\nu}(h)$ as the sum of pheromone values and safety values between the neighboring waypoints $\mu$ and $\nu$ in the $h$th iteration. The initial pheromones on each edge are equal, namely $\sigma_{\mu\nu}(0)=C$ for all $\mu$ and $\nu$. We denote by $s^{*}$ the global best path obtained by the algorithm, corresponding to the minimal-cost trajectory in our UAV optimization problem. Next, for each ant $k\in\{1,2\ldots,m\}$ in the colony, we initialize the pheromone $V_{P0}$, the safety value $\kappa_{0}$, and the heuristic value $\eta$. We also define the evaporation rate $\rho\in(0,1)$ representing the degree to which the guidance factor $\sigma_{\mu\nu}(h)$ decays with iterations. Finally, we define $\mathscr{S}_{\iota}$ as the set of waypoints that ant $\iota$ can pass next, and $p_{\mu\nu}^{\iota}(h)$ as the probability that ant $\iota$ moves from position $\mu$ to $\nu$ in the $h$th iteration. The exact value of $p_{\mu\nu}^{\iota}(h)$ is jointly determined by the guidance factor and heuristic value at the waypoint, as in the following equation,
$p_{\mu\nu}^{\iota}(h)=\begin{cases}\frac{\sigma_{\mu\nu}^{\alpha}(h)\eta_{\mu\nu}{}^{\beta}(h)}{\sum_{r\in\mathscr{S}_{\iota}}\sigma_{\mu r}^{\alpha}(h)\eta_{\mu r}{}^{\beta}(h)}&\nu\in\mathscr{S}_{\iota},\\ \quad\quad\quad 0&\text{ otherwise.}\end{cases}$ (24)
@@ 5.4 ACS-DS: Detailed Steps
1:Positions of all MDs $(x_{j},y_{j},z_{j})$$\rho$; $\eta$; $V_{P0}$; $\kappa_{0}$; terrain surface with no-fly zone $(x^{o}_{n},y^{o}_{n},z^{o}_{n})\in\boldsymbol{O}$
[…]
@@
8:    end for
9:    for each ant $\iota$ do
10:     $(x_{\iota}(h),y_{\iota}(h),z_{\iota}(h))$ = $(x(0),y(0),z(0))$
11:     for each edge $(\mu,\nu)$ do
12:      Choose the next coordinate with probability $p_{\mu\nu}^{\iota}(h)$ by $\sigma$ and $\eta$
13:      while $(x_{\iota}(h+1),y_{\iota}(h+1),z_{\iota}(h+1))\notin\boldsymbol{O}$ do
14:         Output $(x_{\iota}(h+1),y_{\iota}(h+1),z_{\iota}(h+1))$
15:      end while
16:     end for
[…]
18:     for each edge $(\mu,\nu)$ do
19:      Update $V_{P_{\mu\nu}}$ and $\kappa_{\mu\nu}$ by $\rho$
20:      Update $\sigma_{\mu\nu}(h)$ by $V_{P_{\mu\nu}}$ and $\kappa_{\mu\nu}$
21:     end for
22:    end for
23:  end while
24:end for
25:Compute and output $E^{\mathrm{MOV}*}(t)$ by (10).
5.5 ACS-DS: Proof of convergence
@@
$\lim_{h\rightarrow\infty}\sigma_{\mu^{\prime}\nu^{\prime}}\left(h^{*}+h^{\prime}\right)$ $=\max\Bigl\{\sigma_{\mu\nu}\left(h^{*}\right),$ (27)
$\qquad\lim_{h\rightarrow\infty}\left[(1-\rho)^{h^{\prime}}\cdot\sigma_{\mu^{\prime}\nu^{\prime}}\left(h^{*}\right)\right]\Bigr\}$
@@
$\leq\sigma_{\mu\nu}\left(h^{*}\right).$
@@ 5.6 ACS-DS: Complexity analysis
6 Performance Evaluation
We now present numerical results to assess the efficacy and versatility of our proposed solutions.
6.1 Experiment setup
(a) Side view
(b) Top view
Figure 7: Trajectories planned by different mechanisms on a 3D spatial domain. The terminal point shown in Fig. 7 represents the UAV’s final hovering location for sustained communication rather than the physical position of the MD.
We consider a three-dimensional spatial domain with dimensions of $S\times S\times Z$. We randomly generate a continuous surface (no-fly zone) within a three-dimensional spatial domain as shown in Figs. 7(a) and 7(b), and consider that $K$ MDs are placed on the ground conforming to the terrain height, with their positions uniformly and randomly distributed in the space. All MDs are placed at the terrain surface height, and the UAV always operates at an altitude higher than the MD elevations. The values of key system parameters are listed in Table 5.
Table 5: VALUES OF KEY SYSTEM PARAMETERS
@@ 6.1 Experiment setup
$m$ 50 $\mathrm{~kg}$ $a$ 5.28 $\mathrm{~m}/\mathrm{s}^{2}$
@@
$T$ 20–35 min $N_{j}$ 1–5
$f^{\min}$ 1.0 GHz $f^{\max}$ 3.0 GHz
$f_{j}^{\min}$ 0.5 GHz $f_{j}^{\max}$ 2.0 GHz
[…]
$\bar{Q}$ $5\times 10^{7}$ (bit-equivalent) $\tau$ 1.0 s
The UAV flight time $T$ is obtained by cumulatively summing the propulsion energy consumed at each timeslot, where the energy usage is determined from the rotor-speed–dependent motor model already defined in Section 3.2. The endurance is reached once the accumulated consumption equals the battery capacity $B_{c}$, giving the maximum feasible $T$ under a given trajectory and control strategy. Specifically, $T$ is calculated as:
$T=\max\left\{t\in\mathbb{N}:\sum_{\tau=0}^{t}E^{\text{UAV}}(\tau)\leq B_{c}\right\},$ (28)
where $E^{\text{UAV}}(\tau)$ is defined in (20) and includes both movement energy $E^{\mathrm{MOV}}(\tau)$ defined in (10)) and other energy consumption components. Values of key parameters used in the algorithms described earlier in this paper are listed in Table 6. All experimental results presented in this section are based on the average of the 50 independent runs with the weighting factor $\epsilon$ randomly generated within its domain for each run.
Table 6: VALUES OF KEY ALGORITHMIC PARAMETERS
@@
For the task assignment and resource allocation part, we adopt PSO (Algorithm 1), which has been demonstrated effective and robust in our earlier conference paper [24], for all experiments in this section. For the attitude control and trajectory planning modules, we focus on evaluating the performance of our proposed ACS-DS (Algorithm 2) and ATC (Section 5.1). Note that our work addresses a joint UAV control–communication–computation optimization problem, which fundamentally differs from traditional networking-only scheduling tasks. Conventional networking baselines such as Round Robin, Proportional Fair, or queue-based schedulers cannot operate in our setting, as they assume fixed infrastructure and do not interact with UAV flight dynamics, propulsion energy, or real-time orientation-dependent channel variations. These classical methods also do not model UAV physical constraints, motor-level energy consumption, computation frequency control, or end-to-end latency decomposition, all of which are essential components of our optimization problem. Therefore, consistent with the standard practice in the UAV-assisted MEC and cross-layer optimization literature, we adopt representative algorithmic families that can jointly optimize continuous control variables and discrete resource allocation decisions. Specifically, we examine the following ten implementations, including our proposed approach, recently proposed heuristic (e.g., [48, 44]), and RL-based methods [35], and compare their performances in terms of overall operational efficiency cost.
  • 1.
@@
    ACS: The trajectory of the UAV is planned using the classical Ant Colony System (ACS). Task assignment, processing frequencies, transmission powers, and channel allocation follow the optimization procedures described in Section 3. No special mechanisms are applied for attitude control.
  • 2.
    GA-SCA: The UAV trajectory is optimized using a hybrid method combining Genetic Algorithm (GA) and Successive Convex Approximation (SCA), as proposed in [48]. The remaining parameters follow the same strategy as ACS.
  • 3.
    CPS-ACO: The UAV trajectory is planned using the Chaotic-Polarized-Simulated Ant Colony Optimization (CPS-ACO) method, proposed in [44]. Other system parameters are optimized as in ACS.
  • 4.
    TD3: The UAV trajectory is planned using the Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm, as described in [35, 47]. Other configurations are consistent with the baseline setup in ACS.
  • 5.
    ACS-DS: This is an enhanced version of ACS, where UAV trajectory planning incorporates our proposed decoupling and safety value mechanisms, detailed in Algorithm 2. Other parameter settings remain unchanged.
  • 6.
    CPS-ACO+ATC: Based on CPS-ACO, this method integrates the attitude control mechanism described in Section 5.1 to enable more stable UAV dynamics.
  • 7.
    TD3+ATC: Based on TD3, this variant incorporates the attitude control mechanism described in Section 5.1 to improve stability and responsiveness.
  • 8.
    ACS-DS+ATC: This is the full version of our proposed method, which combines the decoupled safe trajectory planning in ACS-DS with the attitude control mechanism in Section 5.1.
6.2 Simulation Environments
The experimental evaluation of the proposed framework is conducted using three complementary simulation environments, each designed to address different aspects of system validation and serving distinct roles in the evaluation pipeline.
(a) Isaac sim deployment simulation
(b) Web-based demonstration
(c) AirSim deployment simulation (AirSimNH)
(d) AirSim deployment simulation (Landscape Mountains with snow)
Figure 8: Simulation environments used for framework evaluation.
  • 1.
    Isaac Sim: This Omniverse/PhysX-based robotics simulation platform provides high-fidelity rigid-body UAV dynamics that accurately model inertia, thrust, and torque characteristics. As shown in Fig. 8(a), the environment incorporates realistic environmental disturbances including wind fields, actuator latency, and platform vibration, enabling high-fidelity physical verification of control algorithms. We utilize Isaac Sim to validate the FEAR-PID controller’s stability under realistic dynamics and to verify that the UAV parameters specified in Table 5 exhibit physically correct behavior. We additionally leverage Isaac Sim’s built-in/official USD environment assets as realistic scene data for representative outdoor scenarios, so that the controller is also inspected under non-trivial geometry and terrain context beyond synthetic setups. This physics-based validation ensures that our control strategies and parameter selections are feasible for real-world deployment.
  • 2.
    Microsoft AirSim: This platform offers photorealistic rendering powered by the Unreal Engine, featuring realistic suburban and mountainous environments that represent the large-scale outdoor deployment scenarios described in our system model (e.g., AirSimNH and LandscapeMountains scenes). As shown in Fig. 8(c) and Fig. 8(d), the environment provides high-fidelity visual representation of complex terrain with realistic lighting and atmospheric effects. The platform provides comprehensive sensor emulation including RGB cameras, depth sensors, semantic segmentation, IMU, GPS, and realistic noise models, and also exposes environment controls such as time-of-day and weather. We employ AirSim to support interactive mission execution with the built-in multirotor dynamics under deployment-scale scenes and configurable conditions, and to generate synchronized FPV/chase-view visualizations for representative runs via ExternalPhysicsEngine (pose-driven) replay, which facilitates consistent presentation and debugging across different methods.
  • 3.
    Lightweight web-based demonstration [^2] Online demo: https://shuaijun-liu.github.io/UAV-Assisted-Fog-Computing-Simulation-Demo: This accessible interactive visualization tool is inspired by AirSim’s visual style. As illustrated in Fig. 8(b), the simplified environment employs streamlined UAV dynamics without full rigid-body physics, yet incorporates dynamic wind effects, simplified control loops, and procedural terrain generation. The web demo serves to qualitatively visualize UAV behavior, trajectory evolution, and controller stability, offering practitioners and researchers an intuitive understanding of the system’s operational characteristics without requiring specialized simulation software.
These three environments complement each other: Isaac Sim validates physical feasibility and control robustness with high-fidelity dynamics and asset-based scenes, AirSim evaluates deployment-scale performance in interactive Unreal Engine environments, and the web demonstration provides intuitive qualitative insights. Together, they form a comprehensive evaluation framework that spans from low-level control validation to high-level mission assessment.
6.3 Module-wise Ablation Study
To evaluate the individual contribution of each proposed component in our holistic framework, we conduct an ablation study by selectively removing one module at a time while keeping the others unchanged. This experimental design allows us to quantify the performance degradation caused by the absence of specific functionalities and to demonstrate the necessity of integrating all components for optimal system performance. We consider the following four key modules:
  • 1.
    FEAR-PID-based Attitude Control (ATC): Enhances UAV stability and reduces energy overhead during flight maneuvers;
  • 2.
    ACS-DS Trajectory Planner: Provides obstacle-aware and energy-efficient path planning with fast convergence. Determining the operational efficiency costs of the movement;
  • 3.
    PSO-based Task and Resource Assignment: Allocates task execution and communication resources effectively. Determining the operational efficiency costs of tasks computation and unloading.;
  • 4.
    Gamma-based Channel Allocation Strategy: Prevents bandwidth monopolization by larger tasks. Determining the operational efficiency costs of transmission.
Each of these modules is removed individually to create a controlled ablation variant of the full framework. The tested ablation variants are summarized in Table 7.
Table 7: Ablation variants with specific module removals.
Variant Name Description
Full All modules enabled (ATC + ACS-DS + PSO + Gamma)
w/o ATC Replaces FEAR-PID with classical PID control
w/o ACS-DS Replaces ACS-DS with TD3 trajectory planner
w/o PSO Uses random task assignment and uniform power allocation
w/o Gamma Uses uniform channel assignment ($P_{j}(t)=1/K$)
Baseline Classical PID + TD3 trajectory +
uniform task/power/channel assignment
For each variant, we use the same network environment and task configuration as in the full model experiments. All experiments are repeated 50 times with randomized task generation to ensure statistical robustness. The evaluation metric is the overall operational efficiency cost, defined as the sum of total task delay and energy consumption, normalized across scenarios.
6.4 Numerical results
6.4.1 Optimal Propeller Parameters